OpenStudy (anonymous):

The Cobb family decided to go roller skating for their reunion. There were a total of 44 people there and all went roller skating. If an adult admission was $5.00 and a child's admission was $3.00 and they paid a total of $168.00, how many children were at the reunion? @reesyboo

OpenStudy (anonymous):

5A+3C=168 A+C=44 5A+3(44-A)=168 2A=168=132 A=18 C=26

OpenStudy (anonymous):

yea i got the same answer as u